OpenStudy (anonymous):

I need help! Find the slope of the line that passes through each pair of points. 1. M(-2,3), N(7,-4) 2. J(-9,3),K(2,1)

OpenStudy (esshotwired):

Do you know point-slope form?

OpenStudy (anonymous):

Huh?

OpenStudy (esshotwired):

point-slope form is what you use for questions like this.

OpenStudy (anonymous):

nope...

OpenStudy (esshotwired):

ponit-slope form is this: y – y1 = m(x – x1) but the ones should be underscored

OpenStudy (esshotwired):

It is just something use use when you know two order pairs.

OpenStudy (anonymous):

my brain hurts from looking at that...

OpenStudy (esshotwired):

haha it really isn't bad you just have to plug the numbers in.

OpenStudy (esshotwired):

Do you know how to find the slope with two ordered pairs?

OpenStudy (anonymous):

^^That's what we're supposed to do!

OpenStudy (anonymous):

So, no.

OpenStudy (esshotwired):

\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] This is the equation you use to find the slope

OpenStudy (anonymous):

Whats the y and the I?

OpenStudy (esshotwired):

Well to find the slope for M(-2,3), N(7,-4) you have to plug it in. the the numbers that are underscored tell which ordered pair it is from. So y2 would be the y value in the 2nd ordered pair (7,-4)

OpenStudy (anonymous):

OMG I DON'T GET THIS AT ALL

OpenStudy (esshotwired):

let me find a website that might help you.

OpenStudy (esshotwired):

http://www.mathwarehouse.com/algebra/linear_equation/slope-of-a-line.php use this one to help you find the slope

OpenStudy (anonymous):

I know how to find the slope with a graph/ just not that

OpenStudy (esshotwired):

This is a good website that explains it all. If you need help with more explanation just let me know. http://www.mathsisfun.com/algebra/line-equation-point-slope.html

OpenStudy (anonymous):

I still don't get it...

OpenStudy (esshotwired):

Try this. Graph the line for 1 and then find the slope that way. When you are done let me know.

OpenStudy (anonymous):

Ugh. This is so difficult. I can't draw a graph.

OpenStudy (anonymous):

I don't have graph paper.

OpenStudy (esshotwired):

The website I just put is graph paper. If you give me a few minutes I will solve them.

OpenStudy (esshotwired):

\[y=-\frac{ 7 }{ 9}x-\frac{ 14 }{ 9 }\] This is number 1

OpenStudy (esshotwired):

\[y=-\frac{ 2 }{ 11 }x+\frac{ 15 }{ 11 }\] This is number 2